What is the probability of being dealt 4 clubs from a deck of 52 cards?
13/52*12/51*11/50*10/49
= 1/4*4/7*11/50*10/49
= 440/68600
= 44/6860
= 11/1715
What is the probability that at least 2 students in a class of 36 have the same birthday?
About 83.2%
The probability that non of the 36 students have the same birthday (not considering
February 28 of the leap year) is given by the following relation:
P(non out of n have same bd) = Π1n-1 [(365-i)/365]
P(non out of 36 have same bd) = (364/365)(363/365)(362/365) ... (331/365)(330/365) =
= 0.167817892.. ≈ 16.8%
So the probability of at least 2 having the same birthday is about 1 - .168 = 0.832 =
83.2%
What is the probability drawing a black king from a pack of cards?
Assuming there are no Joker cards the chance is one in twenty six.
There are fifty two cards in a pack and only two of them are black kings.
What will be the sampling distribution of the mean for a sample size of one?
It will be the same as the distribution of the random variable itself.
What is the probability that you toss a head or a tail?
Assuming a two-sided coin, and that you make the the toss, the probability of tossing a head or a tail is 100%. The probability of tossing a head is 50%. The probability of tossing a tail is 50%.
How many permutations are in the word peak?
The first letter of the perm can be any of the 4, the second any of the remaining 3, the third either of the remaining 2. There are thus 4 x 3 x 2 ie 24 permutations.
What number can be used to describe a certain event?
The number one can be used to describe a certain event.
How many different ways cant the letters deeds be arranged?
They can't be arranged in a million different ways!
What is the probability of rolling an even number if you rolled a six-sided die?
3 out of 6 or 1 out of 2
3/6 or 1/2
Which of the following best explains step 2 of the ORM six step process?
the assessment step involves the application of quantitative or qualitative measures to determine the probability and severity of ill effects potentially resulting from exposure to a hazard
What is the probability of drawing a red king from a deck of 52?
Probability of getting a King is 4/52 = 1/13. Having drawn the King, for it to be a Red King, the probability is 2/4 = 1/2. Hence the conditional probability is (1/13)x(1/2) = 1/26
What is the probability of spinning an even number all four times?
That depends on the numbers you can spin.
What is the probability of getting two prime numbers 1-10?
The probability of getting two prime numbers when two numbers are selected at random and without replacement, from 1 to 10 is 2/15.
Is continuous distribution normal distribution?
le standard normal distribution is a normal distribution who has mean 0 and variance 1
Can a z score be a negative value?
Yes. However, because the distribution is symmetric about 0, some tables give only positive values for z.
What do you call the bell shaped curve that represents a symmetrical distribution of scores?
You call it a bell shaped curved. It may or may not be Gaussian (Normal).
Are probability and random sampling anonymous?
Probability is a branch of mathematics and so is not linked with any individual and so is anonymous. Random sampling may or may not include information that will allow the contributor to be identified. So it may or may not be anonymous.
Why 1.34 cannot be the probability of some event?
1.34 cannot be the probability of an event because the range of possibility is 0 to 1. 0 is an impossible event and 1 is a certain event. There cannot be a possibility higher than certain so 1.34 is not possible.